You have just received notification that you have won the $2 million first prize in the Centennial Lottery. However, the prize will be awarded on your 100th birthday (assuming you’re around to collect), 67 years from now.

  

What is the present value of your windfall if the appropriate discount rate is 9 percent? (Do not round intermediate calculations and round your final answer to 2 decimal places, e.g., 32.16.)
Answer #1

The present value of the windfall is :

The future value as estimated 67 years from now is $2,000,000.

The time period is 67 years

I/Y = 9%

So, the PV is :

FV = $2,000,000

N = 67 Years

I/Y = 9%

So, PV = $6348.1122

= $6348.11( ROUNDED OFF TO TWO DECIMAL PLACES)

This can also be calculated as :

PV = FV/ (1 + r)^n

= $2,000,000/ (1.09)^67

= $2,000,000/ 321.7921

= $6215.19 ( rounded off to two decimal places)